OpenStudy (anonymous):

Determine if y is a factor or a term: 8(x2 + x)y plz help explain to me how to solve this

OpenStudy (anonymous):

anyone plz help me

OpenStudy (gw2011):

8(x^2+x)y = 8y(x^2+x) y is treated just like 8--the only difference is that the y was placed after the parenthesis in the original problem but it means the same thing as the 8. In other words, you multiply (x^2+x) by both the 8 and y.

OpenStudy (anonymous):

okay i get that i think but i dont have to work the problem out i just have to determine if y is a factor or a term

OpenStudy (gw2011):

Okay--therefore, if 8 is a factor, then it makes y a factor also since it is outside the parenthesis.

OpenStudy (anonymous):

okay so than any number outside the parenthesis would be a factor?

OpenStudy (gw2011):

Essentially, if a number or letter is outside the parenthesis it is a factor because it simplifies the problem. For example, (x^2+x) in the problem you show can be further simplified by the following factor: (x^2+x) = x(x+1)

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

You're welcome